OpenStudy (anonymous):

use the slope formula to determine the slope of the line containing the two points. (-1,7) and (-5,-3)

OpenStudy (anonymous):

\(\sf \LARGE \color{maroon}{slope \ = \ \frac{ y _{2} \ - \ y _{1} }{ x _{2} \ - x _{1} }}\)

OpenStudy (anonymous):

do you want the proof?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

in mathematics - or in physics - slope is \(\sf \Large \tan \theta \)

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

but you need to know that in physics , \(\sf \Large \tan \theta \) is called \(\sf \large \color{green}{The \hspace{5pt} Acceleration} \)

OpenStudy (anonymous):

and do you want to know how \(\sf \large \tan \theta \) becomes (y2 - y1)/(x2-x1) ?